I drag a mass m=21 kg in a straight line, along a horizontal surface, a distance D=23 m . I drag it at constant speed v=0.90 m.s−1 in a straight line using a horizontal force. The coefficients of friction are \mu_{s} = 1.2μ
s
​ =1.2 and \mu_{k} = 1.1μ
k
​ =1.1. How much work do I do?
1
Expert's answer
2020-03-09T11:15:09-0400

As the object is in motion then kinetic friction force will act on it and to keep the object in motion we have to apply a constant force which will be equal to the magnitude of kinetic friction force

Normal Force = "m\times g=21\times 10=210"


"f_k=\mu_k\times Normal\\f_k=1.1\times 210=21\times11=231 N"


"\\Work \ done=Force\times displacement \\Work \ done=231\times23=5313J"